OpenStudy (anonymous):

100 people rate a restaurant. 45 of 60 men liked it while 36 of 40 woman did. What is P(Woman|Liked the restaurant)?

OpenStudy (anonymous):

36/40 = 0.9

OpenStudy (anonymous):

can u tell me whose probability we hv to find ???

OpenStudy (anonymous):

It's just P(Woman|Liked the restaurant) or the probability that you choose a woman given that they liked the restaurant

OpenStudy (anonymous):

The correct answer was .56 but I don't know how?

OpenStudy (dumbcow):

aww probability \[P(A |B) = \frac{P(AB)}{P(B)}\] P(AB) is probability of person being woman AND liking restaurant....36/100 meet that criteria P(B) is prob of liking restaurant (45+36)/100 meet that criteria \[P(A|B) = \frac{.36}{.81}\]

OpenStudy (dumbcow):

answer is 0.56 ? that would be true if question was P(Man | like restaurant)

OpenStudy (anonymous):

That's so weird, maybe the question was wrong. Cause I got 36/81 which is .44. Thanks though.

OpenStudy (dumbcow):

welcome...yeah maybe its a typo but im 99% sure answer is 0.44 like you said